McCann Catching, Inc. has 2.00 million shares of stock outstanding. The stock currently sells for $12.90 per share. The firm**THIS QUESTION HAS 4 PARTS TOTAL**

1. What is the cost of equity?

2. What is the percentage of equity used by McCann Catching, Inc.?

3. What is the WACC for McCann Catching, Inc.?

4. What is the NPV of the expansion? (answer in terms of millions, so 1,000,000 would be 1.0000)
Answer #1

Equity Market Value = 2,000,000 * 12.87 = 25,740,000

Debt Market Value = 16,000,000 * 89% = 14,240,000

Toal Capital = 25,740,000 + 14,240,000 = 39,980,000

Cost of Equity = Rf + beta * Market Risk Premium

= 4% + 1.49 * 8%

= 15.92%

WACC = (Cost of Equity * Weight of Equity) + (Cost of Debt after tax * Weight of Debt)

= 15.92% * 25,740,000 / 39,980,000 + 10% * (1-0.38) * 14,240,000 / 39,980,000

= 12.46%

NPV = Present Value of Cash Inflow - Initial Outlay

= Annual Earnigns * PVAF (12.46%, 8 years) - 47,380,000

= 11,000,000 mn * 4.8880637914 - 47,380,000

= 6,396,870